f(x) = 3(x-1) (x+2) / (x-4) (x+2)

a) state vertical asymptotes

b) state the roots:


1
Expert's answer
2021-09-13T16:31:12-0400

"\frac{3(x-1) (x+2)}{ (x-4) (x+2)}=\frac{3(x-1)}{ (x-4)}\\
a) x+4=0=>x=-4\\
x=-4\, \text{is the vertical asymptotes.}\\
b) 3(x-1)=0=>x=1\\
x=1\,\text{is the root of the given function.}"